OpenStudy (anonymous):

how high on a wall does a 12-foot ladder reach if the base of the ladder is 4 feet from the wall? (To the nearest tenth of a foot)

OpenStudy (anonymous):

a^2+b^2=c^2 4^2+b^2=12^2 16+b^2=144 144-16=128 radical 128= 11.313

OpenStudy (anonymous):

Thanks